OpenStudy (anonymous):

May someone please help me solve this problem? I'm factoring a polynomial...I think. \[(x+1)(x^2-x+1)\] I got stuck while attempting to simplify (x-x) and (1-x)

OpenStudy (anonymous):

this is the formula for factoring the sum of two cubes, so before we even start, the answer is \[x^3+1\] but we can work through it step by step

OpenStudy (anonymous):

\[(x+1)(x^2-x+1)\] first multiply all by \(x\) then multiply all by \(1\) you get \[x^3-x^2+x+x^2-x+1\]

OpenStudy (anonymous):

then combine like terms all adds up to zero \(-x^2+x^2=0,-x+x=0\) you are left only with \(x^3+1\)

OpenStudy (anonymous):

I see where I messed up at; thank you for explaining ^-^

OpenStudy (anonymous):

yw